A solenoid is made from a 26-m length of wire of resistivity 2 4x10-8 Ω·m The wire, whose radius is 2.0 mm, is wrapped uniformly onto a plastic tube 4.6 cm in diameter and 1.65 m long Part A Find the emf to which the ends of the wire must be connected to produce a magnetic field of 1.6x10-2 T within the solenoid Express your answer using two significant figures. E- Submit Provide Feedback Next
Answer #1

Given,

L = 26 m ; rho = 2.4 x 10^-8 ; r = 2 mm ; D = 4.6 cm ; l = 1.65 m

A)B = 1.6 x 10^-2 T

We know that

B = u0 N I/L

I = BL/u0 N

N = 1.65/2 x 3.14 x 0.023 =

I = 1.6 x 10^-2 x 1.65/(4 x 3.14 x 10^-7 x 11.42) = 1480.55 A

R = rho L/A

R = 2.4 x 10^-8 x 26/(3.14 x (2 x 10^-3)^2) = 0.049

V = IR = 1480.55 x 0.049 = 72.55 Volts

Hence, V = 72.55 V